# Vevo2: A Unified and Controllable Framework for Speech and Singing Voice Generation

### v2

@@ III-B Speech-Singing Joint Training
To conduct the joint training of speech and singing voice, an essential question lies in how to unify the generation process of speech and singing within a single framework. For the acoustic modeling stage, once the proposed unified content-style tokenizer is obtained, this part becomes relatively straightforward, as both speech and singing voice can be formulated into unified content-style tokens and decoded into Mel spectrograms under the guidance of a timbre reference. However, we observe that the main challenge of speech-singing joint training arises in the content-style modeling stage, where the generation characteristics of speech and singing differ substantially. For instance, speech generation like TTS usually takes only text (and the reference) as input and does not use an explicit control over prosody [tts-book-tanxu], while singing voice generation often demands direct control over melody [singing-expression, diffsinger, tcsinger, spsinger].
@@
Motivated by this observation, we focus on the unified prosody learning and propose two different training strategies to bridge speech and singing voice generation during the content-style modeling stage (Figure 3): (1) Implicit Prosody Learning (IPL): the AR model generates content-style tokens solely from text input, implicitly inferring prosodic information through in-context learning. This strategy is commonly adopted in zero-shot TTS systems [audiolm, seedtts, cosyvoice, vevo, maskgct]. (2) Explicit Prosody Learning (EPL): in addition to the text input, prosody tokens are explicitly provided, allowing the AR model to predict content-style tokens conditioned on both text and prosody information. Notably, during the joint training, we do not restrict IPL to speech data and EPL to singing voice data. Instead, we treat all samples equally: for each training sample, we randomly choose between IPL and EPL with equal probability. We assume that this design better unifies the training patterns for speech and singing voice.
@@
During the acoustic modeling stage, we employ an FM transformer to convert the content-style tokens into Mel spectrograms, conditioned on a timbre reference. This process follows the general paradigm used in speech generation studies [seedmusic, cosyvoice, vevo, glm4-iclr], with the distinction that our approach leverages the proposed unified content-style tokenizer and is jointly trained on both speech and singing voice data. After obtaining the Mel spectrograms, we use a Vocos-based vocoder [vocos], which is also trained on both speech and singing voice, to reconstruct the final waveform.
@@ III-C Multi-Objective Alignment
Intelligibility Reward To align intelligibility, we adopt a recent zero-shot TTS intelligibility preference corpus, INTP [intp], which contains diverse intelligibility preference pairs constructed through comparisons between multiple models as well as manually designed intelligibility perturbations, to train an intelligibility reward model. Specifically, we employ the commonly used Bradley-Terry (BT) approach [bt-rm, instructgpt, RLHF-anthoropic] for reward model training. Given an intelligibility preference dataset $\mathcal{D}=\{(t,a^{w},a^{l})\}$, where $a^{w}$ and $a^{l}$ represent the positive and negative audio outputs with respect to the target text $t$, we parametrize an intelligibility reward model $r_{\phi}(t,a)$, treat the problem as binary classification, and optimize the negative log-likelihood loss:
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})=-\mathbb{E}_{(t,a^{w},a^{l})\sim\mathcal{D}}\left[\log\sigma\left(r_{\phi}(t,a^{w})-r_{\phi}(a,y^{l})\right)\right],$ (5)
@@
Multi-Objective Optimization We apply the GRPO algorithm [grpo] to optimize the proposed multi-objective alignment task. Assuming the number of completions per prompt is $K$, for each prompt, after obtaining the intelligibility rewards ($\bm{r}_{int}$) and prosody similarity rewards ($\bm{r}_{pro}$) for all $K$ completions, we apply group normalization to both rewards separately. The sum of these normalized rewards serves as the final outcome supervision:
@@ III-D Inference-Time Controllability
During the content-style modeling stage, for example, when the prosodic source is not used, the inference paradigm is identical to classic zero-shot TTS [valle, seedtts, vevo, cosyvoice]. As illustrated in the TTS task in Figure 1, the process can be represented as $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$], where the style signals in $\boldsymbol{\bm{\hat{cs}}}$ are prompted by $\boldsymbol{\bm{cs}}$ and generated through the in-context learning capabilities of the pre-trained model. Particularly, when both prosodic source and style reference are utilized as inputs, the inference paradigm takes the form $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]. In this context, for $\boldsymbol{\bm{\hat{cs}}}$, we expect $\boldsymbol{\bm{p}}$ to provide coarse-grained prosody contour signals, while $\boldsymbol{\bm{cs}}$ contributes fine-grained style details (e.g., the vibrato patterns for singing). We will discuss this further through our experiments on style-converted singing voice conversion (Section V-C).
@@
Duration Control Some Non-AR based speech generation models have often featured the ability to control the total duration of generated speech [maskgct, f5tts], while AR-based models are generally considered less capable of controlling output duration [audiolm, seedtts, cosyvoice, vevo, indextts2]. Interestingly, however, we discover that Vevo2 can effectively control the duration of generated results through the use of prosody tokens as input. This capability stems from our tokenizers’ fixed frame rates (6.25 Hz and 12.5 Hz), which maintain a consistent 2:1 ratio between content-style and prosody token sequences during training. Consequently, during inference, we can control the duration of the generated output by applying simple linear scaling transformations to the chromagram features of the prosodic source, thereby modifying the length of prosody tokens. Such control mechanism is detailed in Section V-D.
Pitch Region Control For voice conversion (VC) or singing voice conversion (SVC) tasks, due to potential significant differences in timbre types [vocalset, singingvoice-science] between source and reference speakers, we often need to adjust the pitch region (i.e., apply pitch shift) of the source waveform during inference to achieve higher speaker similarity [svcc-2023, amphion-svc, neucosvc]. In Vevo2, we discover that pitch region control can also be achieved: by shifting the pitch of the source waveform before prosody token extraction, we can effectively control the pitch region of the generated speech or singing voice. Our experiments demonstrate that this approach is effective both when using only the FM stage and when combining AR and FM stages. Such control mechanism is detailed in Section V-E.
@@ IV-A Data and Metrics
Training Data For pre-training, we use Emilia [emilia] as the speech data (101K hours), and follow SingNet’s pipeline [singnet] to prepare 7K hours of singing voice data (source-separated from in-the-wild songs from the Internet). For post-training, we utilize INTP [intp] (which contains 250K preference speech pairs) to train the intelligibility reward model. For the proposed multi-objective alignment task, we employ three types of prosody prompts: (1) speech data: 20K positive samples randomly selected from INTP, (2) singing voice data: 20K samples from M4Singer [m4singer], and (3) instrumental sounds: rendered from MIDIs in M4Singer, covering 16 instruments across five categories: pianos, strings, woodwinds, brass, and folk instruments.
Evaluation Data We consider various settings to construct the evaluation set: (1) For regular speech, we utilize SeedTTS test sets [seedtts], which is widely adopted in the TTS field [maskgct, cosyvoice2, f5tts]. (2) For expressive speech, we employ Genshin-Voice [genshin_voice_dataset], which consists of professional voice acting recordings from the game Genshin Impact. (3) For singing voice, we use GTSinger [gtsinger] and SingStyle111 [singstyle111], which encompass a wide range of singing techniques, styles, and timbre types. (4) Additionally, for two unique tasks: humming-to-singing and instrument-to-singing, we utilize humming [humtrans] and instrumental sound including flute [flute-dataset] and violin [violin-dataset] for the evaluation. All evaluations are conducted using target text in both English and Chinese languages.
Evaluation Metrics For the objective metrics, we evaluate the intelligibility (WER, $\downarrow$), speaker similarity (SIM, $\uparrow$), F0 correlation (FPC, $\uparrow$[svcc-2023, amphion-svc, vevo]. For WER, we employ Whisper-large-v3 [whisper] for English texts, and Paraformer-zh [paraformer, funasr] for Chinese texts. For SIM, we compute the cosine similarity between the WavLM TDNN [wavlm] speaker embeddings of generated samples and the reference speeches. For subjective metrics, we employ Comparative Mean Opinion Score (CMOS, rated from -2 to 2, $\uparrow$) to evaluate naturalness (N-CMOS) and similarity in speaker, prosody, melody, and style (SS-CMOS, PS-CMOS, MS-CMOS, and Style-CMOS). Besides, we use Melody-MOS to evaluate the melody-following ability (rated from 1 to 3, corresponding to “unable to follow”, “roughly following the melody contour”, and “completely following all melody details”). Detailed backgrounds of subjects and definitions of all subjective metrics are provided in the supplementary materials.
@@ IV-B Implementation Details
Prosody Tokenizer and Content-Style Tokenizer This paper processes and trains on speech and singing voice data with a sampling rate of 24 kHz. For chromagram feature extraction, we employ Librosa’s implementation [^4] https://librosa.org/doc/main/generated/librosa.feature.chroma_stft.html with the following parameters: number of FFT points set to 1920, hop length of 480, window size of 1920, and 24 chroma bins. For Whisper hidden features [whisper], we utilize the encoder output from the pre-trained whisper-medium model [^5] https://huggingface.co/openai/whisper-medium. We adopt the VQ-VAE implementations from Amphion [^6] https://github.com/open-mmlab/Amphion/tree/main/models/vc/vevo [amphion, amphion_v0.2], which serves for MaskGCT [maskgct] and Vevo [vevo]. Our only modification is the addition of a CNN-based downsampling layer before the encoder input and a CNN-based upsampling layer after the decoder output. The extracted chromagram features and Whisper features are both at 50 Hz. For the prosody tokenizer, the down(up)-sampling ratio is 8, resulting in a token rate of 6.25 Hz, with a codebook vocabulary size of 512. For the content-style tokenizer, the down(up)-sampling ratio is 4, yielding a token rate of 12.5 Hz, with a codebook vocabulary size of 16,384.
@@
All models in this study use AdamW [adam, adamw] as the optimizer. For the two tokenizers, we use a peak learning rate of 1e-4, with linear warm-up for 10K steps followed by decay over the remaining training period. Both tokenizers are trained for 300K updates. The prosody tokenizer and content-style tokenizer contain 38M and 44M parameters, respectively.
@@
Content-Style Modeling During the content-style modeling stage, we initialize the AR model using Qwen2.5-0.5B [^7] https://huggingface.co/Qwen/Qwen2.5-0.5B [qwen2.5] and expand its original vocabulary to accommodate our proposed prosody tokens and content-style tokens. The total parameter count of the AR model is 509M. For the BPE-based text tokenizer, we adopted the same tokenization strategy as CosyVoice 2, which masks out one-to-many tokens to prevent excessive pronunciation length and mitigate corner cases caused by data sparsity [cosyvoice2]. During pre-training, we use a peak learning rate of 5e-4, with linear warm-up for 32K steps followed by decay over the remaining training period. The model is pre-trained for 500K updates.
@@
Acoustic Modeling During the acoustic modeling stage, we adopted the flow-matching (FM) transformer architecture from Vevo [vevo]. The transformer consists of 16 layers with 16 attention heads and a hidden dimension of 1024. The FM model contains 363M parameters in total. To enhance training efficiency, we incorporated the REPA strategy [repa] and enforced alignment between the hidden output of the transformer’s 5th layer and the W2v-BERT 2.0 features [^8] https://huggingface.co/facebook/w2v-bert-2.0 [w2v-bert, maskgct] during the training process. We use a peak learning rate of 7.5e-5, with linear warm-up for 32K steps followed by decay over the remaining training period. The FM model is trained for 700K updates. To develop a unified vocoder for both speech and singing voice, we fine-tuned Vevo’s speech vocoder [^9] https://huggingface.co/amphion/Vevo [vevo] using our training data (comprising 101K hours of speech and 7K hours of singing voice data). The vocoder, which is based on the Vocos architecture [vocos], contains 255M parameters and underwent fine-tuning for 572K updates.
@@
Post-training During the post-training, for each training iteration with the 20K M4Singer samples, we randomly selected one instrument and rendered the corresponding audio using pretty_midi [^10] https://github.com/craffel/pretty-midi. We use the original text of the sample and extract prosody tokens from these rendered instrumental sounds. We adopt the training framework of TRL [^11] https://huggingface.co/docs/trl/main/grpo_trainer. We use the learning rate of 5e-6, the batch size of 3, the rollout number of 8 and the KL coefficent is 0.1. The model is post-trained for 9K updates.

### v3

@@ III-B Speech-Singing Joint Training
To conduct the joint training of speech and singing voice, an essential question lies in how to unify the generation process of speech and singing within a single framework. For the acoustic modeling stage, once the proposed unified content-style tokenizer is obtained, this part becomes relatively straightforward, as both speech and singing voice can be formulated into unified content-style tokens and decoded into Mel spectrograms under the guidance of a timbre reference. However, we observe that the main challenge of speech-singing joint training arises in the content-style modeling stage, where the generation characteristics of speech and singing differ substantially. For instance, speech generation like TTS usually takes only text (and the reference) as input and does not use an explicit control over prosody [65], while singing voice generation often demands direct control over melody [66, 44, 98, 101].
@@
Motivated by this observation, we focus on the unified prosody learning and propose two different training strategies to bridge speech and singing voice generation during the content-style modeling stage (Figure 3): (1) Implicit Prosody Learning (IPL): the AR model generates content-style tokens solely from text input, implicitly inferring prosodic information through in-context learning. This strategy is commonly adopted in zero-shot TTS systems [6, 1, 21, 95, 71]. (2) Explicit Prosody Learning (EPL): in addition to the text input, prosody tokens are explicitly provided, allowing the AR model to predict content-style tokens conditioned on both text and prosody information. Notably, during the joint training, we do not restrict IPL to speech data and EPL to singing voice data. Instead, we treat all samples equally: for each training sample, we randomly choose between IPL and EPL with equal probability. We assume that this design better unifies the training patterns for speech and singing voice.
@@
During the acoustic modeling stage, we employ an FM transformer to convert the content-style tokens into Mel spectrograms, conditioned on a timbre reference. This process follows the general paradigm used in speech generation studies [4, 21, 95, 85], with the distinction that our approach leverages the proposed unified content-style tokenizer and is jointly trained on both speech and singing voice data. After obtaining the Mel spectrograms, we use a Vocos-based vocoder [62], which is also trained on both speech and singing voice, to reconstruct the final waveform.
@@ III-C Multi-Objective Alignment
Intelligibility Reward To align intelligibility, we adopt a recent zero-shot TTS intelligibility preference corpus, INTP [93], which contains diverse intelligibility preference pairs constructed through comparisons between multiple models as well as manually designed intelligibility perturbations, to train an intelligibility reward model. Specifically, we employ the commonly used Bradley-Terry (BT) approach [7, 53, 5] for reward model training. Given an intelligibility preference dataset $\mathcal{D}=\{(t,a^{w},a^{l})\}$, where $a^{w}$ and $a^{l}$ represent the positive and negative audio outputs with respect to the target text $t$, we parametrize an intelligibility reward model $r_{\phi}(t,a)$, treat the problem as binary classification, and optimize the negative log-likelihood loss:
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})=-\mathbb{E}_{(t,a^{w},a^{l})\sim\mathcal{D}}\left[\log\sigma\left(r_{\phi}(t,a^{w})-r_{\phi}(t,a^{l})\right)\right],$ (5)
@@
Multi-Objective Optimization We apply the GRPO algorithm [60] to optimize the proposed multi-objective alignment task. Assuming the number of completions per prompt is $K$, for each prompt, after obtaining the intelligibility rewards ($\bm{r}_{int}$) and prosody similarity rewards ($\bm{r}_{pro}$) for all $K$ completions, we apply group normalization to both rewards separately. The sum of these normalized rewards serves as the final outcome supervision:
@@ III-D Inference-Time Controllability
During the content-style modeling stage, for example, when the prosodic source is not used, the inference paradigm is identical to classic zero-shot TTS [69, 1, 95, 21]. As illustrated in the TTS task in Figure 1, the process can be represented as $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$], where the style signals in $\boldsymbol{\bm{\hat{cs}}}$ are prompted by $\boldsymbol{\bm{cs}}$ and generated through the in-context learning capabilities of the pre-trained model. Particularly, when both prosodic source and style reference are utilized as inputs, the inference paradigm takes the form $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]. In this context, for $\boldsymbol{\bm{\hat{cs}}}$, we expect $\boldsymbol{\bm{p}}$ to provide coarse-grained prosody contour signals, while $\boldsymbol{\bm{cs}}$ contributes fine-grained style details (e.g., the vibrato patterns for singing). We will discuss this further through our experiments on style-converted singing voice conversion (Section V-C).
@@
Duration Control Some Non-AR based speech generation models have often featured the ability to control the total duration of generated speech [71, 11], while AR-based models are generally considered less capable of controlling output duration [6, 1, 21, 95, 103]. Interestingly, however, we discover that Vevo2 can effectively control the duration of generated results through the use of prosody tokens as input. This capability stems from our tokenizers’ fixed frame rates (6.25 Hz and 12.5 Hz), which maintain a consistent 2:1 ratio between content-style and prosody token sequences during training. Consequently, during inference, we can control the duration of the generated output by applying simple linear scaling transformations to the chromagram features of the prosodic source, thereby modifying the length of prosody tokens. Such control mechanism is detailed in Section V-D.
Pitch Region Control For voice conversion (VC) or singing voice conversion (SVC) tasks, due to potential significant differences in timbre types [73, 64] between source and reference speakers, we often need to adjust the pitch region (i.e., apply pitch shift) of the source waveform during inference to achieve higher speaker similarity [33, 91, 59]. In Vevo2, we discover that pitch region control can also be achieved: by shifting the pitch of the source waveform before prosody token extraction, we can effectively control the pitch region of the generated speech or singing voice. Our experiments demonstrate that this approach is effective both when using only the FM stage and when combining AR and FM stages. Such control mechanism is detailed in Section V-E.
@@ IV-A Data and Metrics
Training Data For pre-training, we use Emilia [30] as the speech data (101K hours), and follow SingNet’s pipeline [27] to prepare 7K hours of singing voice data (source-separated from in-the-wild songs from the Internet). For post-training, we utilize INTP [93] (which contains 250K preference speech pairs) to train the intelligibility reward model. For the proposed multi-objective alignment task, we employ three types of prosody prompts: (1) speech data: 20K positive samples randomly selected from INTP, (2) singing voice data: 20K samples from M4Singer [89], and (3) instrumental sounds: rendered from MIDIs in M4Singer, covering 16 instruments across five categories: pianos, strings, woodwinds, brass, and folk instruments.
Evaluation Data We consider various settings to construct the evaluation set: (1) For regular speech, we utilize SeedTTS test sets [1], which is widely adopted in the TTS field [71, 22, 11]. (2) For expressive speech, we employ Genshin-Voice [61], which consists of professional voice acting recordings from the game Genshin Impact. (3) For singing voice, we use GTSinger [99] and SingStyle111 [16], which encompass a wide range of singing techniques, styles, and timbre types. (4) Additionally, for two unique tasks: humming-to-singing and instrument-to-singing, we utilize humming [45] and instrumental sound including flute [8] and violin [20] for the evaluation. All evaluations are conducted using target text in both English and Chinese languages.
Evaluation Metrics For the objective metrics, we evaluate the intelligibility (WER, $\downarrow$), speaker similarity (SIM, $\uparrow$), F0 correlation (FPC, $\uparrow$[33, 91, 95]. For WER, we employ Whisper-large-v3 [58] for English texts, and Paraformer-zh [26, 25] for Chinese texts. For SIM, we compute the cosine similarity between the WavLM TDNN [10] speaker embeddings of generated samples and the reference speeches. For subjective metrics, we employ Comparative Mean Opinion Score (CMOS, rated from -2 to 2, $\uparrow$) to evaluate naturalness (N-CMOS) and similarity in speaker, prosody, melody, and style (SS-CMOS, PS-CMOS, MS-CMOS, and Style-CMOS). Besides, we use Melody-MOS to evaluate the melody-following ability (rated from 1 to 3, corresponding to “unable to follow”, “roughly following the melody contour”, and “completely following all melody details”). Detailed backgrounds of subjects and definitions of all subjective metrics are provided in the supplementary materials.
@@ IV-B Implementation Details
Prosody Tokenizer and Content-Style Tokenizer This paper processes and trains on speech and singing voice data with a sampling rate of 24 kHz. For chromagram feature extraction, we employ Librosa’s implementation [^4] https://librosa.org/doc/main/generated/librosa.feature.chroma_stft.html with the following parameters: number of FFT points set to 1920, hop length of 480, window size of 1920, and 24 chroma bins. For Whisper hidden features [58], we utilize the encoder output from the pre-trained whisper-medium model [^5] https://huggingface.co/openai/whisper-medium. We adopt the VQ-VAE implementations from Amphion [^6] https://github.com/open-mmlab/Amphion/tree/main/models/vc/vevo [94, 42], which serves for MaskGCT [71] and Vevo [95]. Our only modification is the addition of a CNN-based downsampling layer before the encoder input and a CNN-based upsampling layer after the decoder output. The extracted chromagram features and Whisper features are both at 50 Hz. For the prosody tokenizer, the down(up)-sampling ratio is 8, resulting in a token rate of 6.25 Hz, with a codebook vocabulary size of 512. For the content-style tokenizer, the down(up)-sampling ratio is 4, yielding a token rate of 12.5 Hz, with a codebook vocabulary size of 16,384. The prosody tokenizer and content-style tokenizer contain 38M and 44M parameters, respectively.
@@
All models in this study use AdamW [38, 47] as the optimizer. For the two tokenizers, we use a peak learning rate of 1e-4, with linear warm-up for 10K steps followed by decay over the remaining training period. Both tokenizers are trained on 4 NVIDIA A100 GPUs for 300K updates (about 48 hours).
@@
Content-Style Modeling During the content-style modeling stage, we initialize the AR model using Qwen2.5-0.5B [^7] https://huggingface.co/Qwen/Qwen2.5-0.5B [77] and expand its original vocabulary to accommodate our proposed prosody tokens and content-style tokens. The total parameter count of the AR model is 509M. For the BPE-based text tokenizer, we adopted the same tokenization strategy as CosyVoice 2, which masks out one-to-many tokens to prevent excessive pronunciation length and mitigate corner cases caused by data sparsity [22]. During pre-training, we use a peak learning rate of 5e-4, with linear warm-up for 32K steps followed by decay over the remaining training period. The model is pre-trained on 8 NVIDIA A100 GPUs for 500K updates (about 168 hours).
@@
Acoustic Modeling During the acoustic modeling stage, we adopted the flow-matching (FM) transformer architecture from Vevo [95]. The transformer consists of 16 layers with 16 attention heads and a hidden dimension of 1024. The FM model contains 363M parameters in total. To enhance training efficiency, we incorporated the REPA strategy [82] and enforced alignment between the hidden output of the transformer’s 5th layer and the W2v-BERT 2.0 features [^8] https://huggingface.co/facebook/w2v-bert-2.0 [12, 71] during the training process. We use a peak learning rate of 7.5e-5, with linear warm-up for 32K steps followed by decay over the remaining training period. The FM model is trained on 8 NVIDIA A100 GPUs for 700K updates (about 168 hours). To develop a unified vocoder for both speech and singing voice, we fine-tuned Vevo’s speech vocoder [^9] https://huggingface.co/amphion/Vevo [95] using our training data (comprising 101K hours of speech and 7K hours of singing voice data). The vocoder, which is based on the Vocos architecture [62], contains 255M parameters and underwent fine-tuning on 4 NVIDIA A100 GPUs for 572K updates (about 72 hours).
@@
Post-training During the post-training, for each training iteration with the 20K M4Singer samples, we randomly selected one instrument and rendered the corresponding audio using pretty_midi [^10] https://github.com/craffel/pretty-midi. We use the original text of the sample and extract prosody tokens from these rendered instrumental sounds. We adopt the training framework of TRL [^11] https://huggingface.co/docs/trl/main/grpo_trainer. We use the learning rate of 5e-6, the batch size of 3, the rollout number of 8 and the KL coefficent is 0.1. The model is post-trained on 8 NVIDIA A100 GPUs for 9K updates (about 48 hours).